How far apart must two electrons be if the force between them is 1x 10-12N?


1
Expert's answer
2021-06-24T16:14:19-0400
"F=k\frac{e^2}{r^2}\\(1\cdot10^{-12})=(9\cdot10^{9})\frac{(1.6\cdot10^{-19})^2}{r^2}\\r=1.5\cdot10^{-8}m"